# Rethinking the Learning Paradigm for Facial Expression Recognition

### v2

@@ Abstract
Due to the subjective crowdsourcing annotations and the inherent inter-class similarity of facial expressions, real-world Facial Expression Recognition (FER) datasets usually exhibit ambiguous annotation. To simplify the learning paradigm, most existing methods convert ambiguous annotation results into precise one-hot annotations and train FER models in an end-to-end supervised manner. In this paper, we rethink the existing training paradigm and propose that it is better to use weakly supervised strategies to train FER models with original ambiguous annotation. Specifically, we model FER as a Partial Label Learning (PLL) problem, which allows each training example to be labeled with an ambiguous candidate set. To better utilize representation learning to boost label disambiguation in PLL, we propose to use the Masked Image Modeling (MIM) strategy to learn the feature representation of facial expressions in a self-supervised manner. Then, these feature representations are used to query the probabilities of ambiguous label candidates, which are used as the ground truth in the PLL label disambiguation process. Extensive experiments on RAF-DB, FERPlus, and AffectNet demonstrate that the PLL learning paradigm substantially outperforms the current fully supervised state-of-the-art approaches in FER.
Index Terms:
Facial expression recognition, Partial label learning, Deep learning.
I Introduction
@@
Aiming at analyzing and understanding human emotions, Facial Expression Recognition (FER) [1, 2, 3, 4, 5] is an active field of research in computer vision with practical implications in social robotics, automatic driving, etc. FER leverages these multimedia elements(including images, videos, and audio) to interpret and understand human emotions by analyzing facial features, vocal tones, and other behavioral cues.
@@
In recent years, FER has achieved impressive performance on laboratory-controlled datasets, such as CK+ [6], JAFFE [7] and RaFD [8], which are collected under ideal conditions and annotated by experts. Recently, with the demands of real-world applications, large-scale FER datasets in unconstrained environments, such as FERPlus [9], RAF-DB [10] and AffectNet [11], continue to emerge. However, collecting large-scale datasets with fully precise annotations is labor-intensive and challenging. Consequently, the crowdsourcing approach is commonly utilized for annotation. Due to the subjective perception of annotators and the inherent inter-class similarity of facial expressions, real-world FER datasets usually exhibit ambiguous annotation.
Figure 1: These samples are from the FERPlus dataset, with the corresponding labels in green, which are obtained by voting on the results of the crowdsourcing annotations. This label conversion strategy can introduce artificial noise.
To make it easier to use the crowd-sourcing annotation results, most previous methods convert them into precise one-hot annotations by simply voting [9] or thresholding [10] and train FER models in an end-to-end supervised manner. However, these label conversion strategies do not always yield satisfactory results as shown in fig. 1. The performance of models supervised with such converted labels is inevitably impeded.
In this paper, we rethink the existing training paradigm and propose that it is better to use weakly supervised strategies to train FER models with original ambiguous annotation. Specifically, we model FER as a Partial Label Learning (PLL) problem, where each training instance is associated with a set of candidate labels among which exactly one is true, reducing the overhead of finding the exact label from ambiguous candidates [12].
The major challenge of PLL lies in the label disambiguation. Existing PLL methods [13, 14, 15, 16] state that examples close to each other in the feature space will tend to share identical labels in the label space and their label disambiguation is based on this assumption. However, there is a contradiction here: the direct use of ambiguous annotations inevitably affects the quality of representation learning, which in turn hinders label disambiguation. Errors can accumulate severely, especially with the low-quality initial representations [17].
To mitigate the influence of ambiguous labels on feature learning, we propose to use unsupervised learning to obtain the initial feature representation. Specifically, we propose to use the recently emerged Masked Image Modeling (MIM) to pre-train facial expression representation. MIM is mostly built upon the Vision Transformer (ViT), which suggests that self-supervised visual representations can be done by masking input image parts while requiring the target model to recover the missing contents. Facial expressions are high-level semantic information obtained from the combination of action units in multiple regions of the human face. Vision Transformer can model the relationship between action units with the global self-attention mechanism. MIM pre-training forces the ViT to learn the local facial action units and global facial structures in various expressions [18]. Then the pre-trained ViT is finetuned with an ambiguous label set in the PLL paradigm. Moreover, unlike other tasks, facial expression data exhibit high inter-class similarity, which makes label disambiguation more challenging.
@@
In this paper, we propose to treat the label disambiguation process as querying the confidence of the correlation between each category label and features. Specifically, inspired by DETR [19], we propose a decoder-based label disambiguation module. We leverage learnable label embeddings as queries to obtain confidence in the correlation between features and the corresponding label via the cross-attention module in the Transformer decoder. Finally, we integrate the encoder of the ViT backbone, and the Transformer decoder-based label disambiguation module to build a fully transformer-based facial expression recognizer in the partial label learning paradigm. Our contributions are summarized as follows:
@@
II Related Work
II-A Facial Expression Recognition
Facial Expression Recognition (FER) [20, 21, 11, 22] aims to help computers understand human behavior and even interact with humans. FER can be mainly divided into traditional-based and deep learning-based methods. For the former, it mainly focuses on hand-designed texture-based feature extraction for in-the-lab FER datasets, on which the recognition accuracy is very high.
Since then, most studies tried to tackle the in-the-wild FER task. These large-scale datasets are collected in an unconstrained environment and annotated by crowdsourcing. Due to the subjective perception of annotators and the inherent inter-class similarity of facial expressions, real-world FER datasets usually exhibit ambiguous annotation. Previous methods [23, 24, 25, 26, 27] convert the annotation results into precise one-hot annotations by simply voting [9] or thresholding [10] and train FER models in an end-to-end supervised manner. However, these label conversion strategies do not always yield satisfactory results as shown in fig. 1. The performance of models supervised with such converted labels is inevitably impeded. In recent years, several attempts try to address the ambiguity problem in FER. SCN [20] and RUL [28] relabel noisy samples automatically by learning clean samples to suppress uncertainty from noise labels. DMUE [29] proposes a method of latent label distribution mining and an inter-sample uncertainty assessment, which aims to solve the problem of label ambiguity. However, they either require the noise rate for better noise filtering or rely on sophisticated architectures, which limit their generalization to in-the-wild situations.
Recently, some researchers move from fully-supervised learning (SL) to semi-supervised learning (SSL). Ada-CM [30] proposes an SSL method by adaptive learning confidence margins to utilize all unlabeled samples during training time, which represents a new idea to solve FER tasks by making use of additional unlabeled data. In this paper, we rethink the existing training paradigm and propose that it is better to use weakly supervised strategies to train FER models with original ambiguous annotation. As fig. 1 shows, in-the-wild facial expressions often include compound emotions or even mixture ones. Thus, for these samples with uncertainty, it is natural to have multiple annotated candidates. Consequently, for the first time, a new Partial Label Learning (PLL) paradigm is proposed to address the annotation ambiguity in FER.
II-B Partial Label Learning
Partial Label Learning (PLL) is an important category of weakly supervised learning methods in which each object has multiple candidate labels, but only one of them is the ground-truth label. It is mainly divided into average-based and identification-based methods. The most intuitive solution is average-based methods, which generally consider each label to be equally important during training time, and make predictions by averaging the output of all candidate labels [31, 32]. Others [33, 34] utilize a parametric model to maximize the average scores of candidate labels minus that of non-candidate labels. The identification-based methods normally maximize the output of the most likely candidate labels to disambiguate the ground-truth label. Some typical methods mainly include maximum margin criterion [35], graph-based [36]. With the development of deep learning, more network structures [37, 38] emerge, which rely on the output of the model itself to disambiguate the candidate label sets. Some researchers argue [39, 40, 41] that data points closer in feature space are more likely to be classified as the same label, which relies on a good feature representation. PICO [42] reconciles PLL-based label disambiguation and feature representation based on contrastive learning. However, elaborately designed data augmentation is essential due to its sensitivity to positive sample selection. Another drawback of contrastive learning in FER is that the high inter-class similarity of facial expressions makes it hard to obtain robust feature representations.
II-C Masked Image Modeling
Inspired by Masked Language Modeling (MLM), Masked Image Modeling (MIM) has become a popular self-supervised method. iGPT [43] predicts unknown pixels by operating on known pixel sequences. ViT [44] predicts masked patch prediction by employing a self-supervised approach. Moreover, iBOT [45] achieves excellent performance through siamese networks. However, these methods have the assumption of image semantic consistency. The ability of BEiT [46] to predict discrete tokens depends on the pre-trained model VQVAE [47] that it relies on. As an extension of the ViT work, the main aim of SimMIM [48] is to predict pixels. Unlike the previously mentioned approaches, MAE [49] proposes an encoder-decoder architecture that uses the decoder for the MIM task. Differently, Maskfeat [50] adopts hog descriptor as a prediction target rather than pixels. Currently, the potential benefits of MIM for enhancing PLL performance have not been well studied.
@@ II-C Masked Image Modeling
Figure 2: The overview of our framework. In the left part, we use the pre-trained ViT encoder as our backbone for the feature representation of the 2D image, In the right part, we input the obtained feature from the left part to the transformer decoder with learnable label embedding. After that, we revise $(i-1)_{th}$ confidence and update it to get $i_{th}$ confidence. Finally, the loss is computed between logits and $i_{t}h$ confidence.
III Method
@@
In this section, we introduce a fully transformer-based facial expression recognizer with a new partial label learning paradigm, which mainly consists of a process of robust learning representation and label disambiguation. fig. 2 shows an overview of our framework. Generally, given an in-the-wild FER dataset $\mathcal{D}(\mathcal{X},\mathcal{Y})$, each sample $x_{i}$ is assigned to one of the $K$ deterministic categories, denoted as $y_{i}\in\left\{y_{1},y_{2},...,y_{k}\right\}$,
III-A Partial Label Learning (PLL)
@@
Partial label learning (PLL) aims to predict a true label of the input using a mapping of the classification function. It can tolerate more uncertainty from the label space, and gradually eliminate it in a weakly supervised way during the training. In PLL, each sample in the partially labeled dataset $\tilde{\mathcal{D}}$ is represented by $(x_{i},Y_{i})$, where $Y_{i}$ is a set of candidate labels corresponding to $x_{i}$.
Label set construction. The first step of using PLL is to construct a label set, $Y_{i}=\left\{y_{1},y_{2},...,y_{k}\right\}$, as the input. Given this label set, PLL would then recognize the potential (unseen) true label $y_{i}$ of each sample $x_{i}$ from $Y_{i}$ by training a classifier, which computes the probability of each class in the candidate label set. Here we adopt the method in [51] to determine the potential ambiguous samples. For the ambiguous samples, we consider binary values and assign 1 to the components in $Y_{i}$ that correspond to the top-k predictions and assign 0 to the rest. We normalize each label set $Y_{i}$ to obtain a confidence vector $C_{i}=\left\{c_{1},c_{2},...,c_{k}\right\}$, with all values summing up to 1.
In the training stage, a classifier $\zeta:\mathcal{X}\to\mathcal{C}^{k}$ is trained to update the confidence vector $c_{i}$. We expect the $c_{i}$ to concentrate most of the probability density on the true label $y_{i}$ of the sample $x_{i}$. We use the categorical cross-entropy loss to constrain it as follows:
$\mathcal{L}_{pll}(x_{i},C_{i})=\sum_{j=1}^{k}-C_{i}log(\zeta^{i}(x_{i})),\;s.t.\;\sum_{j=1}^{k}c_{i,j}=1,$ (1)
where $j$ denotes the index of confidence vector $C_{i}$. $\zeta^{i}(\cdot)$ is the softmax layer of the model. In the following, we continue to describe the two key modules in our framework: model pre-training and label disambiguation.
III-B Model Pre-training
@@
Figure 3: We adopt masked image modeling (MIM) for pre-training, which involves predicting the hog descriptor (HOG) with the randomly masked input. After obtaining the pre-trained model, we fine-tune it for the FER task within our partial label learning paradigm.
An important assumption of PLL is that samples close to each other in the sample space are more likely to share an identical label in the label space, which indicates that a good feature representation is important [14, 34, 41]. Based on this assumption, we choose to pre-train the model to achieve a better representation. Since the model is easily misled by ambiguous labels, fully-supervised pre-training may not be a good choice. Therefore, we adopt masked image modeling (MIM) [49], a self-supervised model, which avoids the use of potentially ambiguous labels. The structure of MIM is illustrated in fig. 3. Specifically, in the encoder part, we adopt the encoder of ViT [44], which only acts on the visible and unmasked patches. For the decoder, we adopt hog descriptors as the prediction target of the decoder rather than pixel values [49]. In this way, we expect MIM to learn the local facial action units and global facial structures in various facial expressions.
In FER, focusing on such unit-level information is more meaningful than focusing on pixel-level information. Indeed, pixel-level information can often be distorted by details such as high-frequency nuances, lighting conditions, and fluctuations in contrast. In the pre-training stage, we only keep the encoder for feature extraction and discard the decoder.
III-C Label Disambiguation

### v4

@@ Abstract.
Due to the subjective crowdsourcing annotations and the inherent inter-class similarity of facial expressions, real-world Facial Expression Recognition (FER) datasets usually exhibit ambiguous annotations. To simplify the learning paradigm, most existing methods convert ambiguous annotation results into precise one-hot annotations and train FER models in an end-to-end supervised manner. In this paper, we rethink the existing training paradigm and propose that it is better to use weakly supervised strategies to train FER models with original ambiguous annotations. Specifically, we model FER as a Partial Label Learning (PLL) problem, which allows each training example to be labeled with an ambiguous candidate set. To better utilize representation learning to boost label disambiguation in PLL, we propose to use the Masked Image Modeling (MIM) strategy to learn the feature representation of facial expressions in a self-supervised manner. Then, these feature representations are used to query the probabilities of ambiguous label candidates, which are used as the ground truth in the PLL label disambiguation process. Extensive experiments on RAF-DB, FERPlus, and AffectNet demonstrate that the PLL learning paradigm substantially outperforms the current fully supervised state-of-the-art approaches in FER.
Facial expression recognition, Partial label learning, Deep learning
journal: TOMMcopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Computer vision
1. Introduction
@@
Aiming at analyzing and understanding human emotions, Facial Expression Recognition (FER) is an active field of research in computer vision with practical implications (wang2023turn; wang2024uvmap; wang2023zero; wang2025fully; li2025freeinsert; nie2024t2td) in social robotics, automatic driving, etc.
Figure 1. Random image samples from the FERPlus dataset with potentially noisy labels (colored in red) introduced by simple, voting-based label conversion (barsoum2016training) from crowdsourcing results.
@@
In recent years, FER has achieved impressive performance on laboratory-controlled datasets, such as CK+ (lucey2010extended), JAFFE (lyons1998coding) and RaFD (langner2010presentation), which are collected under ideal conditions and annotated by experts. Recently, with the demands of real-world applications, large-scale FER datasets in unconstrained environments, such as FERPlus (barsoum2016training), RAF-DB (li2017reliable) and AffectNet (DBLP:journals/corr/abs-1708-03985), are required.
Annotating such large-scale datasets by experts is labor-intensive and challenging. Consequently, crowdsourcing becomes increasingly popular. Since the crowdsourced annotations are essentially a rating distribution for each sample, a typical way to leverage them is to convert them into a one-hot label by simple voting (barsoum2016training) or thresholding (li2017reliable) and then use it for a fully-supervised model training. Since crowdsourcing usually yields subjective and ambiguous annotations, the above label conversion may yield ambiguous results as shown in Figure 1, and the subsequent model training inevitably leads to sub-optimal FER models.
Therefore, recent work has started exploring how to learn with noisy labels. A typical solution is to derive a label distribution from the one-hot label and rely on distribution learning (wang2020suppressing; she2021dive). However, this is limited since a potentially correct label may be assigned a relatively low weight, constraining the possibility of overriding the original, noisy label.
To address the above limitation, in this paper, we rethink the existing training paradigm of FER and propose to use weakly-supervised rather than the fully-supervised training. Concretely, we model FER training as a Partial Label Learning (PLL) problem (cour2011learning), where a set of candidate labels is derived from the one-hot ground truth label. This new paradigm avoids relying solely on a single label that is potentially noisy. An important design in our PLL method is to adopt Masked Image Modeling (MIM) (he2022masked) for model pre-training. This is in line with the consensus in PLL research that a higher-quality feature representation leads to a better PLL method (DBLP:journals/corr/abs-2201-08984; DBLP:conf/nips/LiuD12; zhang2016partial; lyu2019gm). The use of the Vision Transformer (ViT) (dosovitskiy2020image) in MIM particularly helps learn the local facial action units and global facial structures in various expressions (he2022masked).
@@
Moreover, in order to address the high inter-class similarity, we propose to treat the label disambiguation process as querying the confidence of the correlation between each category label and features. Specifically, inspired by DETR (DBLP:conf/eccv/CarionMSUKZ20), we propose a decoder-based label disambiguation module. We leverage learnable label embeddings as queries to obtain confidence in the correlation between features and the corresponding label via the cross-attention module in the Transformer decoder. Finally, we integrate the encoder of the ViT backbone, and the Transformer decoder-based label disambiguation module to build a fully transformer-based facial expression recognizer in the partial label learning paradigm.
Our contributions are summarized as follows:
@@
2. Related Work
2.1. Facial Expression Recognition
Facial Expression Recognition (FER) (wang2020suppressing; ng2003sift; DBLP:journals/corr/abs-1708-03985; albanie2018emotion) aims to help computers understand human behavior and interact with humans. Early FER approaches mainly rely on hand-crafted texture or geometric descriptors, which perform well on controlled in-the-lab datasets. With the availability of large-scale in-the-wild datasets annotated via crowdsourcing, recent research has shifted to deep learning–based FER. However, such real-world datasets often contain ambiguous annotations due to subjective perception and the inherent similarity among expressions. Prior methods (DBLP:conf/mm/WangWL19; DBLP:conf/cvpr/YangCY18; vo2020pyramid; farzaneh2020discriminant; gera2021landmark) typically convert the annotation distributions into one-hot labels using majority voting (barsoum2016training) or thresholding (li2017reliable), which inevitably discards ambiguity information and may introduce incorrect supervision, thereby limiting model performance. To mitigate this issue, SCN (wang2020suppressing) and RUL (zhang2021relative) attempt to relabel noisy samples by estimating confidence from clean samples, while DMUE (she2021dive) learns latent label distributions with sample-level uncertainty modeling. Nonetheless, these approaches fundamentally rely on ambiguous one-hot labels and often require estimating noise rates or adopting complex model designs, hindering their robustness in unconstrained scenarios.
More recent FER works further emphasize the necessity of modeling annotation uncertainty. DE-FER (10375788) leverages masked image modeling to obtain discriminative and robust representations under noisy labels, and FSLDL (shin2024noisy) explicitly learns face-specific label distributions to characterize annotation noise. Similarly, MAE-Face (ma2023unified) shows that masked autoencoding benefits affective representation learning, whereas TA-FER (10149390) proposes a transformer-augmented architecture with online label distillation to refine uncertain labels. Beyond noise robustness, open-set FER has also been explored; OPFER (zhang2024open) addresses ambiguous or unseen expressions arising in the wild, and MCR (tan2025mining) adopts label-free consistency regularization to handle noisy supervision without explicit labels. Despite their improvements, these methods ultimately reduce ambiguous annotations to pseudo one-hot targets, leaving the original label uncertainty underutilized. Ada-CM (li2022towards) further explores weak supervision by incorporating unlabeled data through adaptive confidence margins.
However, existing approaches still do not directly preserve or exploit the ambiguous annotation sets produced by crowdsourcing. This gap motivates revisiting the FER training paradigm and exploring methods that can naturally operate on ambiguous annotations without enforcing deterministic supervision.
@@ 2.1. Facial Expression Recognition
Figure 2. The overview of our framework. First, we use the MIM pre-trained ViT encoder as the backbone for representing the features of a 2D image. Second, we input the obtained feature from the left part to the transformer decoder with learnable label embedding. Third, we revise (i1)(i-1)-th confidence and update it to get $i$-th confidence. Finally, the loss is computed between logits and the $i$-th confidence.
2.2. Partial Label Learning
Partial Label Learning (PLL) is a key weakly supervised learning method, where each object has multiple candidate labels, but only one of them is the ground-truth label. It is mainly divided into average-based and identification-based methods. The intuitive, average-based methods generally consider each label to be equally important during training and predict by averaging the output of all candidate labels (hullermeier2006learning; zhang2015solving). Other studies (cour2011learning; zhang2016partial) use parametric models to maximize differences between the average scores of candidate labels and those of non-candidate labels. Identification-based methods normally maximize the output of the most likely candidate labels in order to disambiguate the ground-truth label, mainly by maximum margin criterion (yu2016maximum) or graph-based approaches (wang2021adaptive). Deep learning advances lead to the emergence of network structures (wen2021leveraged; feng2019partial), rely on model output to disambiguate candidate label sets. Some researchers claim (liu2012conditional; zhang2022semi; lyu2019gm; wu2022revisiting) that nearby data points in the feature space tend to share the same label, emphasizing the importance of a robust feature representation. PICO (wang2022pico) reconciles PLL-based label disambiguation and feature representation through contrastive learning. However, elaborately designed data augmentation is essential due to its sensitivity to positive sample selection. Another drawback of contrastive learning in FER is that the high inter-class similarity of facial expressions makes it hard to obtain robust feature representations.
2.3. Masked Image Modeling
Inspired by Masked Language Modeling (MLM), Masked Image Modeling (MIM) becomes a popular self-supervised method. iGPT (chen2020generative) predicts unknown pixels by operating on known pixel sequences. ViT (dosovitskiy2020image) predicts masked patch prediction by employing a self-supervised approach. Moreover, iBOT (zhou2021ibot) achieves excellent performance through siamese networks. However, these methods have the assumption of image semantic consistency. The ability of BEiT (bao2021beit) to predict discrete tokens depends on the pre-trained model VQVAE (van2017neural) that it relies on. As an extension of ViT, the main aim of SimMIM (xie2022simmim) is to predict pixels. Unlike the previously mentioned approaches, MAE (he2022masked) proposes an encoder-decoder architecture that uses the decoder for the MIM task. Differently, Maskfeat (wei2022masked) adopts hog descriptors as the prediction target rather than pixels. Currently, the potential benefits of MIM for enhancing PLL performance have not been well studied.
3. Proposed Methods
@@
In this section, we introduce a fully transformer-based facial expression recognizer with a new partial label learning paradigm, which mainly consists of a process of robust learning representation learning and label disambiguation as shown in Figure 2. Specifically, we explore the benefits of the facial expression representation based on MIM pre-train and the Transformer decoder-based label confidence queries for effective label disambiguation, on which we elaborate in the next sections.
3.1. Problem Formulation
Generally, given an in-the-wild FER dataset $\mathcal{D}(\mathcal{X},\mathcal{Y})$, each sample $x_{i}$ is assigned to one of the $K$ categories, denoted as $y_{i}\in\left\{y_{1},y_{2},...,y_{k}\right\}$ which represents the deterministic category. This simplified form for labels is consistent with the supervised learning setting. However, as shown in Figure 1, most of the FER data is annotated via crowdsourcing, which can introduce labeling noise. Therefore, it is not advisable for most methods to oversimplify the FER task as a supervised learning (SL) problem and apply one-hot labels to all annotated data while ignoring the ambiguity inherent in the label level.
@@ 3.1. Problem Formulation
Partial-label learning (PLL) is a multi-class classification problem. Unlike traditional ones, PLL aims to predict a true label of the input using a mapping of the classification function. However, PLL can tolerate more uncertainty from the label space, and gradually eliminate it in a weakly supervised way during the training. We argue that modeling the FER task as partial label learning is more appropriate. PLL assumes that each input $x_{i}$ must have a true label $y_{i}$ in the corresponding candidate label set $Y_{i}$ as the Eq.(1):
(1) $p(y_{i}\in Y_{i}|x_{i},Y_{i})=1,\forall(x_{i},y_{i})\in\mathcal{X}\times\mathcal{Y},\forall Y_{i}\in\mathcal{S},$
where $\mathcal{S}$ is the collection of all candidate label sets. In addition, different from SL setup, the definition of each sample $x_{i}$ in partially labeled dataset $\tilde{\mathcal{D}}$ is denoted by $(x_{i},Y_{i})$, where $Y_{i}$ is a set of candidate label corresponding to $x_{i}$.
Thus, the aim of partial label learning (PLL) is to recognize the (unseen) true label $y_{i}$ of each sample $x_{i}$ from $Y_{i}=\left\{y_{1},y_{2},...,y_{k}\right\}$ by training a classifier $\zeta:\mathcal{X}\to\mathcal{C}^{k}$, which utilize to compute the probability of each class in the candidate label set. We construct a candidate label set $Y_{i}$ for the sample $x_{i}$ in the FER dataset, where $Y_{i}$ is a set of binary codes. e.g. $y_{k}$ = 1 if the crowdsourcing result contains this category, else $y_{k}=0$. Before training, we normalized each label set $Y_{i}$ to obtain a confidence vector $C_{i}=\left\{c_{1},c_{2},...,c_{k}\right\}$, which represents the probability value of the sample $x_{i}$ in the different categories, it sums to 1. In the training stage, we train a classifier $\zeta:\mathcal{X}\to\mathcal{C}^{k}$ to update the confidence vector $c_{i}$. Ideally, the $c_{i}$ will concentrate most of the probability density on the true label $y_{i}$ of the sample $x_{i}$. We use the categorical cross-entropy loss to constrain it as the Eq.(2) following:
(2) $\hskip-5.69054pt\mathcal{L}_{pll}(x_{i},C_{i})=\sum_{j=1}^{k}-C_{i}log(\zeta^{i}(x_{i})),\;s.t.\;\sum_{j=1}^{k}c_{i,j}=1,$
where $i$ denotes the index of training sample $x_{i}$ and $j$ denotes the index of confidence vector $C_{i}$. And $\zeta^{i}(x_{i})$ is the soft-max layer of the model with the input $x_{i}$. In the next section, we continue to describe the feature representation part of our proposed method.
3.2. PLL with Pre-trained Feature Representation
@@
Figure 3. We adopt Masked Image Modeling (MIM) for pre-training, which involves predicting the hog descriptor (HOG) with the randomly masked input. After obtaining the pre-trained model, we fine-tune it for the FER task within our framework.
In this work, we use MIM pre-train to obtain a better initial feature representation and based on that, use PLL to do label disambiguation. An important assumption of PLL is that samples close to each other in the sample space are more likely to share an identical label in the label space, which indicates that feature robustness is important for label disambiguation. However, as shown in Figure 1, the uncertainty of the FER data leads to ambiguity at the label level, which hinders the feature representation. Based on these observations, as shown in Figure 3, we use the MIM pre-training method to optimize feature representation. Furthermore, since the network is easily misled by the incorrect labels, guiding feature learning by labels alone is difficult for the FER task, which is one of the bottlenecks of most current fully-supervised learning methods. Therefore, we argue that using a self-supervised pre-training manner can help in obtaining a better feature representation by avoiding being misled by incorrect labels.
Inspired by previous work (he2022masked; wei2022masked), we adopt Masked Image Modeling (MIM) to pre-train facial expression representation by predicting hod descriptor. Specifically, in the encoder part, we adopt the same structure with ViT (dosovitskiy2020image) to pre-train, which is only for visible, unmasked patches. This is because the MIM pre-training manner forces the ViT to learn the local facial action units and global facial structures in various expressions. In FER, focusing on the variations in the action units of the facial expression is more meaningful rather than the information embedded in the pixel level. Thus, in the decoder part, we adopt hog descriptors as the prediction target of the decoder rather than pixel values (he2022masked). In addition, predicting pixels is easy to influence by redundant information (high-frequency details, lighting, and contrast variations), which is less important in the FER task. In contrast, hog descriptor (dalal2005histograms) is dedicated to describing the distribution of edge directions within a local subregion. Finally, the output of the decoder can be reshaped to a reconstructed hog prediction of the image. In the training stage, we only keep the encoder for feature extraction and discard the decoder part.
3.3. Label Disambiguation